# Graph-Sequential Alignment and Uniformity: Toward Enhanced Recommendation Systems

### v1

@@ Abstract.
Recommender Systems, Sequential Learning, Graph Neural Network, Alignment and Uniformity
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
@@ 2. Related Work
Alignment and Uniformity in Recommendation Recent advances in unsupervised contrastive representation learning (Wang and Isola, 2020) highlight two crucial properties for high-quality representations: alignment, ensuring positive pairs are close, and uniformity, distributing instances evenly across a hypersphere. Contrastive losses that optimize these properties have been applied in collaborative filtering (CF) (Wang et al., 2022b; Park et al., 2023), GNN-based (Yang et al., 2023; Wang et al., 2024), and sequential recommendation systems (Chung et al., 2023), yielding improved performance over traditional losses like BPR (Rendle et al., 2009), InfoNCE (Gutmann and Hyvärinen, 2010), and CE (Sun et al., 2019). Unlike prior studies, this work investigates the use of these properties for the joint optimization of GNN-based and sequential encoders.

### v2

@@ Abstract.
Recommender Systems; Sequential Learning; Graph Neural Network; Alignment and Uniformity
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715498isbn: 979-8-4007-1331-6/25/04ccs: Information systems Retrieval models and ranking
@@ 2. Related Work
Alignment and Uniformity in Recommendation Recent advances in unsupervised contrastive representation learning (Wang and Isola, 2020) highlight two crucial properties for high-quality representations: alignment, ensuring positive pairs are close, and uniformity, distributing instances evenly across a hypersphere. Contrastive losses that optimize these properties have been applied in collaborative filtering (CF) (Wang et al., 2022b; Park et al., 2023), GNN-based (Yang et al., 2023; Wang et al., 2024; Yang et al., 2024), and sequential recommendation systems (Chung et al., 2023), yielding improved performance over traditional losses like BPR (Rendle et al., 2009), InfoNCE (Gutmann and Hyvärinen, 2010), and CE (Sun et al., 2019). Unlike prior studies, this work investigates the use of these properties for the joint optimization of GNN-based and sequential encoders.
@@ 5. Conclusion
Acknowledgements.
This work is supported in part by NSF under grants III-2106758, and POSE-2346158. Hao Peng is supported by the National Key R&D Program of China through grant 2022YFB3104703, the NSFC through grants 62322202 and 62441612, and the “Pioneer” and “Leading Goose” R&D Program of Zhejiang through grant 2025C02044.